# Semantically-Shifted Incremental Adapter-Tuning is A Continual ViTransformer

### v1

@@ Abstract
Class-incremental learning (CIL) aims to enable models to continuously learn new classes while overcoming catastrophic forgetting. The introduction of pre-trained models has brought new tuning paradigms to CIL. In this paper, we revisit different parameter-efficient tuning (PET) methods within the context of continual learning. We observe that adapter tuning demonstrates superiority over prompt-based methods, even without parameter expansion in each learning session. Motivated by this, we propose incrementally tuning the shared adapter without imposing parameter update constraints, enhancing the learning capacity of the backbone. Additionally, we employ feature sampling from stored prototypes to retrain a unified classifier, further improving its performance. We estimate the semantic shift of old prototypes without access to past samples and update stored prototypes session by session. Our proposed method eliminates model expansion and avoids retaining any image samples. It surpasses previous pre-trained model-based CIL methods and demonstrates remarkable continual learning capabilities. Experimental results on five CIL benchmarks validate the effectiveness of our approach, achieving state-of-the-art (SOTA) performance.
@@ 1 Introduction
In traditional deep learning, the model can access all the data at once and learning is performed on a static dataset. However, in real-life applications, data usually arrives in a stream format with new classes, requiring the model to learn continuously, known as class-incremental learning (CIL). The primary objective of CIL is to enable the model to learn continuously from non-stationary data streams, facilitating adaptation to new classes and mitigating catastrophic forgetting [7]. A number of methods [28, 34, 54] have been devoted to alleviating catastrophic forgetting. Those methods can be mainly divided into replay-based [2, 3, 28], regularization-based [17, 43, 1], and isolation-based methods [30, 23, 24]. However, all these methods assume that models are trained from scratch while ignoring the generalization ability of a strong pre-trained model [5] in the CIL.
Figure 1: Comparison of different parameter-efficient tuning CIL baselines on CIFAR100 dataset. Left: The relationship between the average accuracy of the incremental sessions and the number of tunable parameters. Right: The average performance of old classes and new classes for each PET method.
@@
Pre-trained vision transformer models [5] have demonstrated excellent performance on various vision tasks. Recently, it has been explored in the field of CIL and continues to receive considerable attention [50, 38, 37, 45]. Due to the powerful representation capabilities of pre-trained models, CIL methods based on pre-trained models achieve significant performance improvements compared to traditional SOTA methods which are trained from scratch. CIL with a pre-trained model typically fixes the pre-trained model to retain the generalizability and adds a few additional training parameters such as adapter [4], prompt [15] and SSF [22], which is referred to as parameter-efficient tuning (PET).
@@
Inspired by language-based intelligence, current research in CIL is primarily focused on the prompt-based method [37, 31, 52]. Typically, these approaches require the construction of a pool of task-specific prompts during the training phase which increases storage overhead. Additionally, selecting prompts during the testing stage incurs additional computational costs. Other PET methods as well as fully fine-tuning are still in exploration in the context of CIL. Recently, SLCA [45] proposes fine-tuning the entire ViT and classifier incrementally with different learning rates. However, fine-tuning the entire pre-trained model requires substantial computational resources. In addition, Adam [50] initially explores the application of other PET methods in CIL using first-session adaptation and branch fusion. Training in the first stage and subsequently freezing the model can reduce training time but result in lower accuracy for subsequent new classes. Our linear probing results reveal that the first-session adaptation is insufficient when there is a significant domain discrepancy between downstream data and the pre-trained model.
@@
In this paper, we first revisit different PET methods within the CIL paradigm. As shown in Fig. 1, we observe that adapter tuning [4] is a better continual learner than prompt-tuning [15] and SSF-tuning [22]. When progressively fine-tuning the prompt and SSF parameters, the forgetting of old classes is catastrophic. In comparison, adapter tuning effectively balances learning new classes and maintaining performance in old classes. Unlike prompt-based methods, which require constructing a prompt pool, adapter tuning avoids catastrophic forgetting even sharing the same parameters across learning sessions. Additionally, the adapter balances the number of tuning parameters and model performance compared to fully fine-tuning. Moreover, unlike previous methods that use feature distillation loss to restrict changes in shared parameters as part of overall loss, we analyze that tuning with constraints hinders continual learning from the perspective of parameter sensitivity. Therefore, we train the adapter and task-specific classifier without parameter regularization in each session, allowing for greater plasticity in learning new classes.
@@
As we only train the local classifier in each learning session, we propose to adopt a new classifier retraining method [54, 45, 32] to further improve the CIL performance. First, we implicitly compute the semantic shift [42] of previous prototypes which leverages the semantic shift of current task samples to estimate the change of old classes. Then, we sample several features according to the updated prototypes to retrain the classifier which is more effective than previous methods. The advantages of our proposed method can be summarized as follows: 1) Fine-tuning adapters significantly reduces training costs and improves learning efficiency; 2) We do not need to retain any image samples; 3) The accuracy for new classes is relatively high which verifies the continual learning capacity of the model.
@@
In summary, our proposed learning framework has the following main contributions: (1) Different from various devotion into the prompt-based methods for CIL, we discover that incrementally tuning adapter is a better continual learner even without constructing an adapter-pool; (2) After each session adaptation with local classifier, we propose to retrain a unified classifier with the semantic shift compensated prototypes which can further improve the performance; (3) Extensive experimental results on five CIL benchmarks demonstrate the superiority of the proposed simple but effective methods which achieves the SOTA.
@@ 2 Related Work
2.1 Class-incremental Learning
Class-incremental learning requires the model to be continuously updated with new class instances while retaining old knowledge [49]. Traditional CIL methods can be categorized into replay-based [2, 3, 28], regularization-based [17, 43, 40, 53], and parameter isolation-based methods [30, 23, 24]. Replay-based methods involve retaining or generating samples of previous classes and incorporating them into the current training phase. These methods often employ strategies for sample selection or sample generation to effectively replay past information. Regularization-based methods add constraints or penalties in the learning process which limit the update of the parameters that are important for old classes. Isolation-based methods aim to isolate and update task-specific parameters. By focusing on updating only a subset of parameters, these methods can mitigate catastrophic forgetting. To expand the representative capacity of a model without compromising its existing knowledge, methods for expanding the network have been proposed [41, 34, 48]. These methods dynamically extend the feature extraction network, combined with the replay-based method, achieving dramatic performance improvements.
@@ 2.1 Class-incremental Learning
2.2 Parameter-Efficient Tuning
Parameter-Efficient Tuning can be considered as a transfer learning method. It refers to not performing full fine-tuning on a pre-trained model, instead inserting and fine-tuning specific sub-modules within the network. This approach is initially demonstrated to have effective transfer learning results in NLP [13, 19, 20, 14]. Recently, similar approaches have been applied to vision transformer models as well. AdaptFormer [4] inserts lightweight modules after the MLP layers in the attention module and has been found to outperform full fine-tuning on action recognition benchmarks. Another PET approach SSF [22] surprisingly outperforms other methods in certain tasks even with a smaller number of parameters. Inspired by the prompt approach used in the language model, VPT [15] applies it to visual models and achieves impressive results across various downstream tasks while only introducing a small number of additional parameters. Furthermore, the prompt-based method has also been used in vision-language models [27, 52, 51, 46] to improve performance on various downstream tasks.
2.3 Continual Learning on a Pre-trained Model
The aforementioned CIL methods all involve training the model from scratch, while CIL with pre-trained model [39, 52, 50, 35] has gained much attention due to its strong feature representation ability. L2P [52] utilizes the pre-trained model and learns a set of extra prompts dynamically to guide the model to solve corresponding tasks. DualPrompt [37] proposes to learn of two mutually unrelated prompt spaces: the general prompt and the expert prompt. It encodes task-invariant instructions and task-specific instructions, respectively. CODAPrompt [31] introduces a decomposed attention-based continual learning prompting method, which offers a larger learning capacity than existing prompt-based methods [52, 37]. SLCA [45] explores the fine-tuning paradigm of the pre-trained models, setting different learning rates for backbone and classifiers, and gains excellent performance. Adam [50] proposes to construct the classifier by merging the embeddings of a pre-trained model and an adapted downstream model. LAE [8] proposes a unified framework that calibrates the adaptation speed of tuning modules and ensembles PET modules to accomplish predictions.
3 Methodology
3.1 Preliminary
Class-incremental learning formulation: We first introduce the definition of CIL. Consider a neural network $\mathcal{M}_{\theta}=f_{\theta_{cls}}(\mathcal{F}_{\theta_{bne}}(\cdot))$ with trainable parameters $\theta=\{\theta_{bne},\theta_{cls}\}$. $\mathcal{F}_{\theta_{bne}}$ represents the feature extraction backbone which extracts features from input images and $f_{\theta_{cls}}$ stands for the classification layer that projects feature representations to class predictions. In CIL setting, $\mathcal{M}_{\theta}$ needs to learn a series of sessions from training data $D_{t}=\{(x^{t}_{1},y^{t}_{1}),(x^{t}_{2},y^{t}_{2}),...\},t=1,...,T$ and satisfy the condition $Y(i)\cap Y(j)=\emptyset,i\neq j$, where $Y(i)$ represent the label set in session $i$. The goal of $\mathcal{M}_{\theta}$ is to perform well on test sets that contain all the classes learned denoted as $\mathcal{Y}=Y(1)\cup...Y(t)$ after $t$-$th$ session.
@@ 3.1 Preliminary
$out=x_{i}+s\cdot\sigma(x_{i}*W_{down})*W_{up},\vspace{-2mm}$ (1)
@@ 3.2 Adapter-tuning without parameter constraints
Implementation of adapter-based baselines: During incremental training sessions, only adapter and classifier layers are updated, and the pre-trained ViT model is frozen. As the cosine classifier has shown great success in CIL, we follow ALICE [26] to use the cosine classifier with a margin. The margin hyper-parameter could also be used as a balance factor to decide the learning and retaining. The training loss can be formulated as follows:
@@
where $cos\theta_{j}^{i}=\frac{w_{i}*f_{j}}{||w_{i}||*||f_{j}||}$, $N^{t}$ denotes the number of training samples of the current session, $s$ and $m$ represent the scale factor and margin factor, respectively.
@@
First, we elaborate on why incrementally tuning the adapter is better in the context of CIL. By utilizing the residual structure, the adapter can retain the generalization capabilities from the pre-trained model while adapting to new tasks. The incremental tuning of the adapter exhibits a cumulative learning capability, where the representational capacity of the adapter is further enhanced as the learning sessions progress. In contrast, both SSF and prompt tuning have limitations when it comes to handling CIL. These methods suffer from overfitting to the current distribution. When the shared parameters excessively overfit each current task, the model gradually loses its generalization ability which is harmful for training a unified model for CIL. Then, we try to utilize KD loss to implicitly limit parameter updates and adjust the weighting factor. As shown in Fig. 3, the results demonstrate that unconstrained training is more beneficial for new-classes learning and improving overall performance. Based on this observation, we propose our proposition from the perspective of parameter sensitivity.
@@
Proof: Given the parameter set $\theta=\{\theta_{1},\theta_{2},...\theta_{N}\}$ and training set $D_{t}=(X_{t},Y_{t})$ in $t$-$th$ session, the definition of parameter sensitivity [9, 47] is defined as

### v2

# Continual Adapter Tuning with Semantic Shift Compensation for Class-Incremental Learning

@@ Abstract
Class-incremental learning (CIL) aims to enable models to incrementally learn new classes while overcoming catastrophic forgetting of old classes. The introduction of pre-trained models has brought new tuning paradigms to CIL. In this paper, we first revisit different parameter-efficient tuning (PET) methods within the context of continual learning. It has been observed that continuous adapter tuning with local classifier training demonstrates superiority over prompt-based methods, even without parameter expansion and complex fusion operations during each learning session. Inspired by this, we propose incrementally tuning the shared adapter without imposing parameter update constraints, enhancing the learning capacity of the backbone. We estimate the semantic shift of old prototypes without access to past image samples and update stored prototypes session by session. Additionally, we employ feature sampling from stored prototypes to retrain a unified classifier, further improving its performance. Our proposed method eliminates model expansion and avoids retaining any image samples. It surpasses previous pre-trained model-based CIL methods and demonstrates remarkable continual learning capabilities. Experimental results on several CIL benchmarks validate the effectiveness of our proposed method, achieving state-of-the-art (SOTA) performance.
Index Terms:
Class-incremental Learning, Adapter-Tuning, Semantic Shift Estimation, Vision Transformer
@@ 1 Introduction
In traditional deep learning, the model can access all the data at once and learning is performed on a static dataset. However, in real-life applications, data usually arrives in a streaming format with new classes, requiring the model to learn continuously, known as class-incremental learning (CIL). The primary objective of CIL is to enable the model to learn continuously from non-stationary data streams, facilitating adaptation to new classes and mitigating catastrophic forgetting [1] of past learned classes. Traditional CIL methods [2, 3, 4] have been devoted to alleviating catastrophic forgetting. Those methods can be mainly divided into replay-based [5, 6, 2], regularization-based [7, 8, 9], and isolation-based methods [10, 11, 12]. However, all these methods assume that models are trained from scratch while ignoring the generalization ability of a strong pre-trained model [13, 14] in the CIL.
@@
Recently, pre-trained vision transformer models [13] have demonstrated excellent performance on various vision tasks. It has also been explored in the field of CIL and continues to receive considerable attention [15, 16, 17, 18, 19, 20, 21]. Due to the powerful representation capabilities of pre-trained models, CIL methods based on pre-trained models achieve significant performance improvements compared to traditional SOTA methods which are trained from scratch. CIL with a pre-trained model typically fixes the pre-trained model to retain the generalizability and adds a few additional training parameters such as adapter [22], prompt [23] and SSF [24], which is referred to as parameter-efficient tuning (PET). Simple but effective PET methods [25, 23, 22, 26] can significantly boost the performance of pre-trained models in downstream tasks but also in continual learning scenarios.
Figure 1: Comparison of different parameter-efficient tuning CIL baselines on CIFAR100 dataset. Left: The relationship between the average accuracy of the incremental sessions and the number of tunable parameters. Right: The average performance of old classes and new classes for each PET method. This figure show that Adapter tuning performs best and is more stable across datasets compared to other PET methods.
@@
Inspired by language-based intelligence, current research in CIL is primarily focused on the prompt-based method such as L2P[27], DualPrompt[17], and CODAprompt [28]. Typically, these approaches require the construction of a pool of task-specific prompts during the training phase which increases storage overhead. Additionally, selecting prompts during the testing stage incurs additional computational costs. Other PET methods as well as fully fine-tuning are still in exploration in the context of CIL. Recently, SLCA [18] proposes fine-tuning the entire ViT and classifier incrementally with different learning rates. However, fine-tuning the entire pre-trained model requires substantial computational resources. In addition, APER [15] initially explores the application of other PET methods in CIL using first-session adaptation and branch fusion. Training in the first stage and subsequently freezing the model [15, 29] can reduce training time but result in lower accuracy for subsequent new classes. Our linear probing results reveal that the first-session adaptation is insufficient when there is a significant domain discrepancy between downstream data and the pre-trained model. EASE [19] propose an expandable adapter structure that yields substantial performance improvements over ADAM [15]. However, it requires continuous expansion of the model architecture and extra storage to save the adapters.
@@
In this paper, we first revisit different PET methods within the CIL paradigm. As shown in Fig. 1, we observe that adapter tuning [22] is a better continual learner than prompt-tuning [23] and SSF-tuning [24]. When progressively fine-tuning the prompt and SSF parameters, the forgetting of old classes is catastrophic. In comparison, adapter tuning effectively balances learning new classes and maintaining performance in old classes. Unlike prompt-based methods, which require constructing a prompt pool, adapter tuning avoids catastrophic forgetting even sharing the same parameters across learning sessions. Additionally, the adapter balances the number of tuning parameters and model performance compared to fully fine-tuning. Moreover, unlike previous methods that use feature distillation loss to restrict changes in shared parameters as part of overall loss, we analyze that tuning with constraints hinders continual learning from the perspective of parameter sensitivity. Therefore, we train the adapter and task-specific classifier without parameter regularization in each session, allowing for greater plasticity in learning new classes. In addition, we find that continuously fine-tuning the adapter is also effective in other CIL settings, such as few-shot class-incremental learning (FSCIL) [30] and hierarchical class-incremental learning (HCIL) [31].
@@
Furthermore, as we only train the local classifier in each learning session, we propose to adopt a new classifier retraining method [4, 18, 32] to further improve the CIL performance. First, we implicitly compute the semantic shift of previous prototypes by leveraging the shift of current task prototypes. Different from SDC[33] which estimates the change of old prototypes by calculating the shift of current samples between different learning sessions, we estimate the shift by calculating the change of prototypes between different tasks, thereby reducing computational costs. Then, we sample multiple features according to the updated prototypes to retrain the classifier which is more effective than previous methods without semantic shift estimation. The advantages of our proposed method can be summarized as follows: 1) Fine-tuning adapters significantly reduces training costs and improves learning efficiency compared to training the whole model; 2) We do not need to retain any image samples; 3) The accuracy for new classes is relatively high which verifies the continual learning capacity of the model. Extensive experimental results on seven CIL benchmarks demonstrate the superiority of the proposed simple but effective methods which achieves the SOTA; We also conduct experiments on FSCIL and HCIL setting and gain excellent performance as well.
@@
In summary, our proposed learning framework has the following main contributions: (1) Different from various devotion into the prompt-based methods for CIL, we discover that incrementally tuning adapter is a better continual learner even without constructing an adapter-pool compared with other PET methods; (2) After each session adaptation with local classifier, we propose to retrain a unified classifier with the semantic shift compensated prototypes which can further improve the CIL performance; (3) We apply our method to FSCIL and HCIL settings, demonstrating superior performance compared to other approaches in these contexts. Compared to our CVPR 2024 version [34], in the contribution (2), we propose a more efficient semantic shift estimation method. Furthermore, in our experiments, we extend the experiments on fine-grained datasets to further verify the effectiveness of our method. Moreover, we conduct experiments on other incremental settings.
The rest of this article is organized as follows. We first summarize the existing related work on CIL in Section 2 and provide the preliminaries of the setting and method in Section 3. Then, we give a detailed description of the proposed method in Section 4. To verify the effectiveness of our proposed method, experiments and ablation studies are given in Section 5. Finally, we conclude and discuss the future work in Section 6.
@@ 2 Related Work
2.1 Traditional Class-incremental Learning
Class-incremental learning requires the model to be continuously updated with new class instances while effectively retaining old knowledge [35]. Traditional CIL methods typically train the model from scratch, primarily focusing on mitigating the forgetting of old classes. To alleviate the forgetting, various data-centric approaches have been proposed, including storing past samples [5, 6, 2, 36, 4, 37] or synthesizing image samples with generative models [38, 39, 40]. In addition to data-centric methods, model-centric methods have also been explored. Model-centric methods include expanding the model parameters [41, 3, 42], restricting parameter updates [6, 7, 8, 43], and parameter isolation [10, 11, 12]. Moreover, beyond model-centric and data-centric perspectives, a growing number of methods emphasize the continuous fine-tuning of models with minimal forgetting, proposing various approaches at the algorithm level. For example, knowledge distillation [44, 45, 46, 47], balancing batch normalization [48], feature space shift [49, 50]. These general model-agnostic CIL methods can also be effectively applied to the continuous fine-tuning process of pre-trained models to alleviate forgetting.
2.2 Parameter-Efficient Tuning
Parameter-Efficient Tuning can be considered as a transfer learning method. It refers to not performing full fine-tuning on a pre-trained model, instead inserting and fine-tuning specific sub-modules within the network. This fine-tuning paradigm does not modify the existing pre-trained model parameters and have effective transfer learning results in NLP [51, 52, 53, 25]. Recently, similar approaches have been applied to vision transformer models as well. AdaptFormer [22] inserts lightweight modules after the MLP layers in the attention module and has been found to outperform full fine-tuning on action recognition benchmarks. Another PET approach SSF [24] surprisingly outperforms other methods in certain tasks even with a smaller number of parameters. Inspired by the prompt approach used in the language model, VPT [23] applies it to visual models and achieves impressive results across various downstream tasks while only introducing a small number of additional parameters. Furthermore, the prompt-based method has also been used in vision-language models [14, 27, 54, 55] to improve performance on various downstream tasks.
2.3 Continual Learning on a Pre-trained Model
The aforementioned CIL methods all involve training the model from scratch, while CIL with pre-trained model [56, 27, 15, 57, 58] has gained much attention due to its strong feature representation ability. L2P [27] utilizes the pre-trained model and learns a set of extra prompts dynamically to guide the model to solve corresponding tasks. DualPrompt [17] proposes to learn of two mutually unrelated prompt spaces: the general prompt and the expert prompt. It encodes task-invariant instructions and task-specific instructions, respectively. CODAPrompt [28] introduces a decomposed attention-based continual learning prompting method, which offers a larger learning capacity than existing prompt-based methods [27, 17]. SLCA [18] explores the fine-tuning paradigm of the pre-trained models, setting different learning rates for backbone and classifiers, and gains excellent performance. ADAM [15] proposes to construct the classifier by merging the embedding of a pre-trained model and an adapted downstream model. LAE [59] proposes a unified framework that calibrates the adaptation speed of tuning modules and ensembles PET modules to accomplish predictions. EASE [19] proposes an expandable adapter ensemble approach, integrating it with a pre-trained model to achieve superior CIL performance. However, this method requires additional model size to save the trained adapters. Several methods [57, 60, 61] have been proposed to enhance the incremental performance of the model by leveraging multi-modal information of the pre-trained vision-language model.
2.4 Few-shot Class-incremental Learning and Hierarchical Class-incremental Learning
In addition to the regular CIL setting, two particular CIL settings have garnered significant attention in recent years: few-shot class-incremental learning (FSCIL) and hierarchical class-incremental learning (HCIL). TOPIC [30] first proposes FSCIL and CEC [62] which decouples the encoder and classifier is a widely used baseline. FACT [63] and ALICE [64] employ pseudo-class data augmentation and consider the compatibility for future classes, which achieve significant performance improvements. Limit [65] and SPPR [66] leverage the base-stage data to construct few-shot settings, and then optimize the model to better adapt to the subsequent few-shot learning sessions. SAVC [67] summarizes and builds upon previous work and leverages supervised contrastive learning and pseudo-class construction techniques, which achieve remarkable results. HIL extends the concept of CIL by taking into account the hierarchical relationships between classes. IIRC [68] evaluates the performance of CIL methods in the context of HIL setting. Furthermore, other HIL settings have also been proposed in recent years [69, 31, 43]. Since both FSCIL and HIL are extensions of CIL, we evaluate our method on these two extra settings to validate the generalizability.
@@ 2.4 Few-shot Class-incremental Learning and Hierarchical Class-incremental Learning
3 Preliminary
Class-incremental learning formulation: We first introduce the definition of CIL. Consider a neural network $\mathcal{M}_{\theta}=f_{\theta_{cls}}(\mathcal{F}_{\theta_{bne}}(\cdot))$ with trainable parameters $\theta=\{\theta_{bne},\theta_{cls}\}$. $\mathcal{F}_{\theta_{bne}}$ represents the feature extraction backbone which extracts features from input images and $f_{\theta_{cls}}$ stands for the classification layer that projects feature representations to class predictions. In CIL setting, $\mathcal{M}_{\theta}$ needs to learn a series of sessions from training data $D_{t}=\{(x^{t}_{1},y^{t}_{1}),(x^{t}_{2},y^{t}_{2}),...\},t=1,...,T$ and satisfy the condition $Y(i)\cap Y(j)=\emptyset,i\neq j$, where $Y(i)$ represent the label set in session $i$. The goal of $\mathcal{M}_{\theta}$ is to perform well on test sets that contain all the classes learned denoted as $\mathcal{Y}=Y(1)\cup...Y(t)$ after $t$-thth session.
@@ 3 Preliminary
$out=x_{i}+s\cdot\sigma(x_{i}*W_{down})*W_{up},$ (1)
@@ 4.1 Adapter-tuning without parameter constraints
Implementation of adapter-based baselines: During incremental training sessions, only adapter and classifier layers are updated, and the pre-trained ViT model is frozen. As the cosine classifier has shown great success in CIL, we follow ALICE [64] to use the cosine classifier with a margin. The margin hyper-parameter could also be used as a balance factor to decide between learning new classes and retaining old knowledge. The training loss can be formulated as follows:
@@
[…]
@@
First, we elaborate on why incrementally tuning the adapter is better in the context of CIL. By utilizing the residual structure, the adapter can retain the generalization capabilities from the pre-trained model while adapting to new tasks. The incremental tuning of the adapter exhibits a cumulative learning capability, where the representational capacity of the adapter is further enhanced as the learning sessions progress. In contrast, both SSF and prompt tuning have limitations when it comes to handling CIL. These methods suffer from overfitting to the current distribution. When the shared parameters excessively overfit each current task, the model gradually loses its generalization ability which is harmful for training a unified model for CIL. Then, we try to utilize KD loss to implicitly limit parameter updates and adjust the weighting factor. As shown in Fig. 3, the results demonstrate that unconstrained training is more beneficial for new classes learning and improving overall performance. Based on this observation, we propose our proposition from the perspective of parameter sensitivity.
@@
Proof: Given the parameter set $\theta=\{\theta_{1},\theta_{2},...\theta_{N}\}$ and training set $D_{t}=(X_{t},Y_{t})$ in $t$-thth session, the definition of parameter sensitivity [74, 70] is defined as
@@
Algorithm 1 Semantic Shift Estimation of Prototypes
Input: Current session training data $D_{t}$, Old Class Prototypes $P_{o}$, Old Model $M_{o}$, New Model $M_{n}$.
Output: Updated old class prototypes $P_{o}^{n}$ with semantic shift estimation.
1: Calculate the prototypes of new classes $P_{n}^{o}$ and $P_{n}^{n}$ on both the old model $M_{o}$ and the new model $M_{n}$ using the dataset $D_{t}$.
2: Calculate the gap $\varphi_{i}^{t-1\to t}$ of the new class prototype on the old model and the new model according to Eq. 8.
3:for $i=1$ to $O_{n}$ do
4:  for $j=1$ to $N_{n}$ do
5:   Calculate the cosine similarity $\alpha_{ij}$ of old class prototype $P_{i}$ and new class prototype $P_{j}$.
6:  end for
7:end for
8:for $i=1$ to $O_{n}$ do
9:  Calculate the semantic shift $\widetilde{\Delta}_{i}^{t-1\to t}$ of old class $i$ according to Eq. 10.
10:  Add the semantic shift $\widetilde{\Delta}_{i}^{t-1\to t}$ to the old class prototype and get $P_{o}^{n}$.
11:end for